compression. D) isobaric compression. 5) The average molecular kinetic energy of a gas can be determined by knowing A) only the number of molecules in the gas. B) only the volume of the gas. C) only the pressure of the gas. D) only the temperature of the gas. E) All of the above quantities must be known to determine the average molecular kinetic energy sion. 10) When a fixed amount of ideal gas goes through an adiabatic expansion, A) its internal (thermal) energy does not change. B) the gas does no work. C) no heat enters or leaves the gas. D) its temperature cannot change. E) its pressure must increase. 6) If we double the root-mean-square speed (thermal speed) of the molecules of a gas, then A) its temperature must increase by a factor of 4. B) its temperature must increase by a factor of 2. its temperature must increase by a factor of D) its pressure must increase by a factor of 2. E) its pressure must increase by a factor of 4. Snrino 2019
Answer #1

5)(D)only the temperature of the gas

Reason:

KE = 3/2 k T

6)(A)its temperature must increase by a factor of 4

Reason:

v(rms) = sqrt (3 R T/M)

if v(rms) doubled, temperature should be 4 times.

10)(A)its internal (thermal) energy does not change.

Reason:

U = Q - W